OpenStudy (anonymous):

on a farm there are three kinds of animals: pigs, cows and horses. 25% are pigs and 40% are cow. if there are 20 pigs how many horses are there?

OpenStudy (anonymous):

if 20 are pigs and 25% are pig then the total number of animals is 20*4 = 80 so 40% of 80 is 32.

OpenStudy (anonymous):

why did you times the 20 by 4?

OpenStudy (anonymous):

think about how 4 parts of 25% (or 1/4) make up the whole so 4 *25% = 100% in this case the value of 25% is 20 pigs to find out how many farm animals there are total you have to multiply the value you know (20 pigs) by 4 --> meaning there are 80 animals on the farm from there you find out how many horses there are

OpenStudy (anonymous):

*cows not horses

OpenStudy (anonymous):

ok i get it...so there are 32 cows on the farm

OpenStudy (anonymous):

so there are 48 horses?

OpenStudy (anonymous):

well if there are 80 total animals and we know 20 are pig and 32 are cows you just have to set up 80-(20 +32) = # of horses